determine the set of values of k for which the line 2y=x+k does not intersect the curve y=x^2-4x+7
Answer by stanbon(75887)

determine the set of values of k for which the line 2y=x+k does not intersect the curve y=x^2-4x+7
-----
Since the quadratic opens up you cannot get above it for all values of "x".
So go below it.
---
The minimum of the quadratic occurs when x = -b/(2a) = 4/2 = 2
f(2) = 4-8+7 = 3
---
Minimum at (2,3)
----
Find the line equation where 2y=x+k passing thru (2,3)
---
2*3 = 2+k
k = 4
-----
Equation:
2y= x+4
y = (1/2)x+2
----
Answer: k < 4
======================
Cheers,
Stan H.
===============